Suppose that the Market for Cigarette is facing the Demand function Q = 20 – 2P and Supply function Q = 10.5 + 0.5P: a) What is the effect on the Equilibrium Price and Quantity when Government imposes a 7% of tax as percent of equilibrium price on each unit of Cigarette produced? [5 marks] b) What is the price elasticity of demand at equilibrium after tax and comment on the answer? [5 marks]


1
Expert's answer
2021-04-19T18:48:05-0400

Solution:

a.). At equilibrium: D = S

20 – 2P = 10.5P + 0.5P

20 – 10.5 = 0.5P + 2P

9.5 = 2.5P

P = 3.8

Equilibrium price = 3.8

Substitute for Q:

20 – 2P = 20 – 2(3.8) = 20 – 7.6 = 12.4

Q = 12.4 units

Equilibrium quantity = 12.4 units


If the government imposes a 7% of tax as a percent of equilibrium price on each unit of Cigarette produced, then the equilibrium price will increase and the equilibrium quantity will decrease.


Calculate the new equilibrium price and quantity:

D: Q = 20 – 2P

Make P the subject:

2P = 20 – Q

P = 10 – 0.5Q

S: Q = 10.5 + 0.5P

Q – 10.5 = 0.5P

P = 2Q – 21

New supply function after tax:

P = 2Q – 21 + 0.07P

P – 0.07P = 2Q – 21

0.93P = 2Q – 21

P = "\frac{2Q}{0.93}" – 22.581


At equilibrium:

D = S

10 – 0.5Q = "\frac{2Q}{0.93}" – 22.581


10 + 22.581 ="\frac{2Q}{0.93}" + 0.5Q


32.581 = 2.151Q + 0.5Q

32.581 = 2.651Q


Q = "\frac{32.581}{2.651}" = 12.29


Q = 12.29 units

New equilibrium quantity = 12.29 units

Substitute for Price:

P = "\frac{2Q}{0.93}" – 22.581 = "\frac{2(12.29)}{0.93}" – 22.581 = 26.43 – 22.581 = 3.849


P = 3.849

New equilibrium price = 3.849

 

b.).  Price elasticity of demand at equilibrium after-tax and comment on the answer?


Elasticity of demand (PED) = "\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


% change in qty demanded = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \times 100 = \frac{12.29 -12.4}{(12.29+12.4)/2 } \times 100"


"\frac{-0.11}{12.345} \times 100 = -0.0089\times 100 = -0.89\%"


% change in price ="\frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 } \times 100 = \frac{3.849 -3.8}{(3.849+3.8)/2 } \times 100"


"\frac{0.049}{3.8245} \times 100 = 0.0128\times 100 = 1.28\%"


Elasticity of demand (PED) = "\frac{-0.89\%}{1.28\%} = -0.695"


Elasticity of demand (PED) = 0.695 or 0.70

The elasticity of demand is less than 1, which means that the market for cigarettes is price inelastic and a normal good. Demand for the product does not change significantly after a price increase due to taxes.